OpenStudy (anonymous):

An object moves in uniform circular motion at 25 m/s and takes 1.0 second to go a quarter circle. What is the radius of the circle ? helpppppp me

OpenStudy (anonymous):

If it takes 1 second to go a quarter circle, it takes 4 seconds to go the whole circle. The circumference of the circle is the distance traveled, so that's \(2\pi r\). \[v=\frac{2\pi r}{T}\] where v = 25 m/s and T = 4.0 s Solve for r

OpenStudy (anonymous):

heres the thing i know that it takes 4 second to go the whole circle and i know thats the formula but idk how to plug them i keep getting wrong numbers :(

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

put up what you have so I can see what you're doing

OpenStudy (anonymous):

4x25 =100 for a whole circle

OpenStudy (anonymous):

idk im lost

OpenStudy (anonymous):

(2xpix4)/1 ??

OpenStudy (anonymous):

v is the velocity, so plug in 25 for v T is the time, that's 4 You should have this.\[25=\frac{ 2\pi r }{ 4 }\]

OpenStudy (anonymous):

r is what you're solving for

OpenStudy (anonymous):

so multiply both sides by 4 to get \(100=2\pi r\)

OpenStudy (anonymous):

Now divide both sides by \(2\pi\) and let me know what you get

OpenStudy (anonymous):

i got 0.062

OpenStudy (anonymous):

oh no i mean 157.07

OpenStudy (anonymous):

I got 15.9

OpenStudy (anonymous):

see 0.06 is one of my answers but 157.07 is not and that what i had gotton before you come to help me

OpenStudy (anonymous):

Is that one of your answers?

OpenStudy (anonymous):

how you got 15.9 ?

OpenStudy (anonymous):

I have 16 in my answers i just need know how you got it so i would do it myself next time please

OpenStudy (anonymous):

multiply 2π first then divide 100 by that 100/(2π) = 100/6.28 = 15.9

OpenStudy (anonymous):

yeah it would be 16 to 2 sig figs

OpenStudy (anonymous):

oh thats explains it yeah i put them all together in the calc and thats why it gave me 157 THANK you so MUCH

OpenStudy (anonymous):

you're welcome.